# 0.5 eV QCD Axion Cosmology

### v48

@@ 9 Conclusions
where $\alpha=e^{2}/(\hbar c)=1/(137.035~999~18~(2))$ (P. J. Mohr et al., 2025) [1.5 parts per billion] is the best present estimate for the dimensionless electromagnetic “fine-structure” coupling constant, and where we have used, in the second line, the relation for the QCD axion decay constant, $f_{A}=\sqrt{\chi_{\rm QCD}}/(m_{A}c^{2})$ (S. Weinberg, 1978; F. Wilczek, 1978), together with the values for the axion rest-mass energy, $m_{A}c^{2}$, given in Eq. (6), and the QCD topological susceptibility, $\chi_{\rm QCD}$, given by Eq. (3) (See Appendix B for the estimate of the uncertainty in the value of $f_{A}$). To estimate the value of $C_{A\gamma}$, we use its relation with the light-quark mass ratio, $m_{u}/m_{d}=0.485~(19)$ (Z. Fodor et al., 2016), and with the electromagnetic-to-strong axial anomaly ratio, $E/N=8/3$, (R. D. Peccei, 1989; M. Gorghetto & G. Villadoro, 2019),

### v49

@@ IX Conclusions
where $\alpha=e^{2}/(\hbar c)=1/(137.035~999~18~(2))$ (P. J. Mohr et al., 2025) [1.5 parts per billion] is the best present estimate for the dimensionless electromagnetic “fine-structure” coupling constant, and where we have used, in the second line, the relation for the QCD axion decay constant, $f_{A}=\sqrt{\chi_{\rm QCD}}/(m_{A}c^{2})$ (S. Weinberg, 1978; F. Wilczek, 1978), together with the values for the axion rest-mass energy, $m_{A}c^{2}$, given in Eq. (9), and the QCD topological susceptibility, $\chi_{\rm QCD}$, given by Eq. (VII) (See Appendix B for the estimate of the uncertainty in the value of $f_{A}$). To estimate the value of $C_{A\gamma}$, we use its relation with the light-quark mass ratio, $m_{u}/m_{d}=0.485~(19)$ (Z. Fodor et al., 2016), and with the electromagnetic-to-strong axial anomaly ratio, $E/N=8/3$, (R. D. Peccei, 1989; M. Gorghetto & G. Villadoro, 2019),